OpenStudy (anonymous):

find the area of a rhombus wiht a perimeter of 40 and an angle of 60

OpenStudy (anonymous):

area=(side length)^2 sin theta

OpenStudy (anonymous):

A=10^2 sin 60=50sqrt{3}

OpenStudy (anonymous):

Perimeter is 40, so side is 40/4=10 You know the height, 10sin(60)=5sqrt(3) Thus you can find the area!

OpenStudy (anonymous):

amogh..there is the square of the side length...no?

OpenStudy (anonymous):

Its basically side * sine of side.

OpenStudy (anonymous):

Which seems to be a square.

OpenStudy (anonymous):

its side^2 sin (angle)

OpenStudy (anonymous):

I wrote the same, Its side* height =side*side*sine of the side

OpenStudy (anonymous):

then it should be 50sqrt(3)

OpenStudy (anonymous):

sine of angle, iguess :S

OpenStudy (anonymous):

then how do i find the area

OpenStudy (anonymous):

uzma gave the answer to that,

OpenStudy (anonymous):

i posted it earlier

OpenStudy (anonymous):

Softball go to your previous question

OpenStudy (anonymous):

A trig solution: \[100\text{ Sin}\left[\frac{60}{180}\pi \right]=86.6025\]

OpenStudy (anonymous):

uzma got it first.